Calculate the path length in centimetres which would give a percent transmittance of 4.57% of a colored dye with a molar absorptivity of 1.38×104 L mol-1 cm-1 at a wavelength of 798 nm if the concentration of the dye is 3.93×10-5 M.
Answer #1

% of absorbance = 100 - 4.57 = 95.43 %

and

absorbance = 0.9543

we know,

absorbance =  molar absorptivity * C * pathlength

or

0.9543 =  1.38×10^4 * 3.93×10^-5 * pathlength

or

pathlength = 1.76 cm (answer)